OpenStudy (anonymous):

An apple falls out of a tree from a height of 2.3 m. What is the impact speed of the apple? A. 45.1 m/s B. 6.71 m/s C. 3.12 m/s D. 9.11 m/s

OpenStudy (anonymous):

First you have to know how long the fall took d=1/2 gt^2 or t^2=2 d/g=2*3.4/9.8

OpenStudy (irishboy123):

you can use v^2 = u^2 + 2 a x, one of the 3 commonly quoted equations of motion, to solve direct. v = final velocity u = initial velocity = 0 a = gravity x = distance